# King’s College London, U.K

### 14 Mar 2024 v3

@@ Abstract
We demonstrate that Assembly Theory, pathway complexity, the assembly index, and the assembly number are subsumed and constitute a weak version of algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity reliant on an approximation method based upon statistical compression, their results obtained due to the use of methods strictly equivalent to the LZ family of compression algorithms used in compressing algorithms such as ZIP, GZIP, or JPEG. Such popular algorithms have been shown to empirically reproduce the results of AT that were reported before in successful application to separating organic from non-organic molecules and in the context of the study of selection and evolution. We prove the connections and full equivalence of Assembly Theory to Shannon Entropy and statistical compression, and AT’s disconnection as a statistical approach from causality. We demonstrate that formulating a traditional statistically compressed description of molecules, or the theory underlying it, does not imply an explanation or quantification of biases in generative (physical or biological) processes, including those brought about by selection and evolution, when lacking in logical consistency and empirical evidence. We argue that in their basic arguments, the authors of AT conflate how objects may assemble with causal directionality, and conclude that Assembly Theory does not explain selection or evolution beyond known and previously established connections, some of which are reviewed.
@@ 1 Introduction
Assembly Theory (AT) has recently garnered significant attention in some academic circles and in the scientific media. Responding to how open-ended forms can emerge from matter without a blueprint, AT purports to explain and quantify selection and evolution. The central claim of Assembly Theory advanced in [1, 2] is that objects with a high assembly index “are very unlikely to form abiotically”. While this has been contested in [3], we will demonstrate the connections of Assembly Theory and its assembly index to well-established theories that the authors claim are different and unrelated to Assembly Theory [4]. We will show that the claims advanced by its authors that Assembly Theory unifies life and biology with physics [2, 5] rely on a circular argument and the use of a popular compression algorithm, with no empirical or logical support to establish deeper connections to selection and evolution not already made before [6, 7]. We demonstrate how Assembly Theory is a weak version of algorithmic complexity (a.k.a. Kolmogorov-Solomonoff-Chaitin complexity) [8, 9, 10, 11], denoted by $\mathbf{K}$, that its proposed measure is an LZ compression scheme and asymptotically bounded by (Shannon) Entropy in a fashion that is in no way dissimilar to previous information and graph-theoretic approaches to chemical complexity. We conclude that Assembly Theory cannot offer a different or better explanation of selection or evolution than the connections already established.
@@ 2.1 The assembly Index is a compression algorithm of the LZ family
This limitation is attributable to the authors’ exclusive consideration of computer programs adhering to the form of ‘Template Program A’:
@@
Any approximation to $\mathbf{K}$ that accounts for identical repetitions, including all known lossless statistical compression algorithms, can achieve equivalent or superior results, as demonstrated in the Suplementary Information and [12]. This alignment with ‘Template Program A’ effectively highlights the association of Assembly Theory with well-established principles of compression and coding theory, thereby refuting the initial claim of its authors to be presenting a unique methodology. Moreover, the authors’ suggestion that their index may be generalised as a universally applicable algorithm for any object (including text) based on the LZ family of compression algorithms further underscores the deep disconnection between Assembly Theory and their authors’ suggestion to reinvent text compression (e.g. ZIP or GZIP) and the current state of the art in the field of statistical and non-statistical compression beyond LZW.
It is worth noting that although there may be minor variations in the implementation details of the assembly index and the assembly number of which the authors themselves have proposed significantly different versions, we demonstrate that they are all qualitatively and quantitatively equivalent to the LZ compression algorithms introduced in the 1970s, such as LZ77/LZ78. See Supplementary Information A.
@@ 2.2 A compression algorithm is a mechanical procedure that corresponds to a physical process
The authors of AT argue that traditional data compression algorithms are an overly abstract process unsuited for modelling the construction (or assembly) of objects. This view overlooks the practical and mechanistic nature of compression algorithms, particularly those in the LZ family. Since their introduction in 1977, LZ algorithms have been effectively used in detection, identification, clustering, and classification across various fields, including biology, chemistry, and medicine  [23, 18, 24], and to approximate algorithmic complexity [10, 25].
@@
On the contrary, one of the distinctive features of algorithmic information theory (AIT), algorithmic complexity being one of its indexes, is that its importance and pervasiveness in mathematics, theoretical computer science, and complexity science are, in fact, established because of the invariance of its results with respect to the model of computation, whether abstract or physically implemented. More specifically, such statements from the authors of AT indicate a fundamental misunderstanding of equivalence classes closed under reductions (in this case, computability classes), which are basic and pivotal concepts in computer science, complexity theory, and mathematics in general. It is evidence in support of this point that none of the formal proofs in the Supplementary Information section of this paper made any mention of any Turing machine.
@@
In a published paper [4], the authors attempted to prove the computability of their assembly index and to distance themselves from algorithmic complexity in this way. Their algorithm, categorised as Template Program A, is trivially computable and requires no proof of computability but also all other resource-bounded approximations to $\mathbf{K}$ are computable, including LZW that has been used for decades for this purpose [10]
@@ 2.3 Conflating object assembly process and causation directionality
The authors assume and present the sequential nature of their algorithm as an advantage. Assembly Theory claims to be able to extract causal knowledge in the form of chains of cause and effect from its assembly pathways. The authors’ central assumption is that each step of the basic Template Program A constitutes a cause-and-effect chain corresponding to how an object may have been physically assembled. This is no different from, and is indeed a restricted version of, a pushdown automaton capable of instantiating grammar compressor (which the authors rename as ‘assembly pathway’) and works exactly like an LZ algorithms.
@@
We have demonstrated that the assembly index requires a finite automaton to instantiate a restricted grammar (see Supplementary Information A), a version of a specific-purpose machine, just as any other resource-bounded approximation to $\mathbf{K}$ would need to be instantiated (e.g. calculated, even by hand). In this case, the size of their formula or the size of the implementation of their computer program is the size of the special-purpose finite automaton, which is common and of fixed size for their calculations (which allows them to be discounted from the final length, just as it is from $\mathbf{K}$ or resource-bounded approximations).
@@ 2.4 Lack of control experiments and absence of supporting empirical data beyond current domain knowledge
Turing’s motivation was to explore what could be automated in the causal mechanisation of operations (see Fig. 1), for example, stepwise by hand using paper and pencil. The shortest among all the computer programs of this type is an upper bound approximation to $\mathbf{K}$. In other words, $\mathbf{K}$ cannot be longer than the length of this diagram. The concept of pathway complexity and copy number instantiated by the assembly index, is a restricted version of statistical nature of a Turing machine or finite automaton.
Figure 1: A: The authors of AT have suggested that $\mathbf{K}$ would be proven to be contained in Assembly TheoryT [32]. This Venn diagram shows how Assembly Theory (AT) is connected and subsumed within algorithmic complexity or ($\mathbf{K}$) and within the group of statistical compression as proven in this paper (see Sup. Inf.). B: Causal transition graph of a Turing machine with number 3019 (in Wolfram’s enumeration scheme [33]) with an empty initial condition found by using a computable method (e.g. CTM, [34]) to explain how the block-patterned string 111000111000 was assembled step-by-step based on the principles of $\mathbf{K}$ describing the state, memory, and output of the process as a fully causal mechanistic explanation. A Turing machine is simply a procedural algorithm and any algorithm can be represented by a Turing machine. By definition, this is a mechanistic process, and as physical as anything else, not an ‘abstract’ or ‘unrealisable’ process.
2.5 Statistical difference between a random versus a nonrandom selection of objects does not unify physics and biology
Central to Assembly Theory (AT) and the public claims made through their University Press Releases [5] is the connection that the authors of AT want to make to selection and evolution, maintaining that AT unifies physics and biology and explains and quantifies selection and evolution, both Darwinian and abiotic [2]. To demonstrate this connection, by assuming selectivity in the combination of linear strings $(P)$, the authors compare two schemes for combining linear strings $Q$, one random versus a non-random selection. This experiment yields observations of differences between the two $(R)$, and the authors conclude that selectivity $(S)$ exists in molecular assembly, and therefore that AT explains it.
@@
Yet, assuming selectivity in the combination of strings and proceeding to say that a selection algorithm is expected to differ from a random one makes for a circular argument.
Furthermore, the conclusion $(S)$ reaffirms the initial assumption $(P)$, resulting in a circular argument.
@@
These circular arguments, lacking a foundation in chemistry, biology, or empirical evidence, lead the authors to propose the concept of “assembly time”. The suggestion is a time-scale separation between molecule production (assembly time) and discovery time, aiming to unify physics and biology. Time, however, was already fundamental in evolutionary theory. Assembly Theory revisits existing complexity science concepts as described in the timeline provided in Fig. 2.
Figure 2: A timeline of results in complexity science relevant to the claims and results from Assembly Theory. Assembly Theory renames several concepts such as dictionary trees ‘assembly pathways’, algorithmic probability and their reduction of combinatorial space. The assembly index and number are equivalent to LZ lossless compression algorithms with the full details provided in the Sup. Inf.
@@ 3 Conclusions
If Assembly Theory has some elements that appear sound, or to the extent that its index displays some discriminatory power [36, 1], it is because the assembly index is a resource-bounded approximation to $\mathbf{K}$ instantiated by a compression algorithm.
@@
We have shown how the assembly index results can be explained in the context of a dictionary-based compression algorithm that belongs to the LZ family. The most popular examples used by the authors of Assembly Theory, such as ABRACADABRA and BANANA are traditionally used to teach LZ77/LZ78 compression in Computer Science courses at university level. Results obtained by Assembly Theory to date have been reported previously in the context of algorithmic information theory [18, 6]. The concepts, results, index, and threshold, as reported in [1, 36, 4, 2], follow directly from earlier work on algorithmic complexity and data compression.
Based on their high and low copy number feature driving pathway complexity and their assembly number, an object with a low assembly index will have low Entropy and high LZ compressibility, and an object with a high assembly index will have high Entropy and low LZ compressibility with respect to the size of the object in direct proportion and driven by the same copy-counting mechanism (on which every popular statistical compression is based upon).
In contrast, algorithmic (Kolmogorov) complexity and other computable approximations [17, 21] potentially have more discriminatory power because they may tell apart cases in which high assembly index values with low copy number, high Entropy, and low LZ compressibility are causally compressible by exploiting short rewriting describing the object beyond statistical identical copies.
@@
Algorithmic complexity has been shown to be profoundly connected to causality [37, 38, 21, 39]. The claim that it can only be applied to strings is incorrect, as it has been applied before to images [40], networks [7, 41, 42, 43], vectors, matrices, and tensors in multiple dimensions [44, 45, 46]; and even chemical structures [18].
@@
One positive aspect of Assembly Theory is that it may offer a graph-like representation approach (albeit more convoluted than necessary) to approximate LZ compression in the specific context of molecular complexity—given that dictionary trees have been used for pedagogic purposes in computer science for decades—but potentially making it more accessible to a wider and less technical community. However, it is inaccurate and a bad practice to ignore, omit or imply the absence of previous work (e.g. [18, 6, 47]), belittle and devalue the theories and ideas upon which their index is fully dependable and connected with (Shannon Entropy and Kolmogorov complexity), and make disproportionate public claims related to the overstated contributions of AT (even calling it a ‘theory of everything’).
We agree with the authors on the importance of understanding the origin and mechanisms of the emergence of an open-ended generation of novelty [48, 39]. However, these efforts by the authors are undermined by hyperbolic claims. The suggestions made by Assembly Theory that people may resonate with highlight the working hypothesis that information theory, algorithmic complexity, and algorithmic probability underpin the processes and building blocks of physical and living systems on which all these methods and frameworks have been heavily inspired upon [49, 50], playing an important role in chemical processes [18], nature and biology as suggested before [18, 51, 6, 40].
@@ A.3.1 Assembly index defines an encoding scheme
    The following results show that the assembly index $c_{\Gamma}\left(y\right)$ is a parameter that can be employed to encode an object $y$. Theorem 5 demonstrates that any arbitrarily chosen encoding of the set of vertices (in any arbitrary ordering) of a minimum rooted assembly space is sufficient to encode the highest-assembly-index object. In other words, the sheer collection of objects in a minimum rooted assembly subspace carries enough information for allowing one to always pick the object with the highest assembly index value, even though the ordering of the objects in this collection can be completely arbitrary in the first place. This demonstrates that Assembly Theory is one of the possible methods to encode objects.

### 19 Mar 2024 v4

@@ Abstract
We demonstrate that Assembly Theory (AT) is reliant on a method based upon statistical compression that belongs to the LZ family of compression algorithms, family which comprises popular algorithms used in ZIP, GZIP, or JPEG. Such popular algorithms have been shown to empirically reproduce the results of AT that were reported before in successful applications to separating organic from non-organic molecules and in the context of the study of selection and evolution. In particular, assembly index is a LZ compression scheme whose statistical compressibility is bounded by entropy or other traditional LZ compression schemes, such as LZ77, LZ78, or LZW. In addition, we demonstrate that AT, its pathway complexity, the assembly index, and the assembly number define compression schemes that are subsumed into the theory of algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity. Due to AT’s current lack of logical consistency in defining causality for non-stochastic processes and lack of the empirical evidence that it outperforms other complexity measures found in the literature, we conclude that the assembly index and the assembly number do not lead to an explanation or quantification of biases in generative (physical or biological) processes, including those brought about by (abiotic or Darwinian) selection and evolution.
@@ 1 Introduction
Assembly Theory (AT) has recently garnered significant attention in some academic circles and in the scientific media. Responding to how open-ended forms can emerge from matter without a blueprint, AT purports to explain and quantify selection and evolution. The central claim of Assembly Theory advanced in [1, 2] is that objects with a high assembly index “are very unlikely to form abiotically”. While this has been contested in [3], we will demonstrate the connections of Assembly Theory and its assembly index to well-established theories that the authors claim are different and unrelated to Assembly Theory [4]. We will show that the claims advanced by its authors that Assembly Theory unifies life and biology with physics [2, 5] rely on a circular argument and the use of a popular compression algorithm, with no empirical or logical support to establish deeper connections to selection and evolution not already made before [6, 7]. We demonstrate how Assembly Theory is an approximation to algorithmic complexity (Kolmogorov-Solomonoff-Chaitin complexity) [8, 9, 10, 11], denoted by $\mathbf{K}$, based on LZ compression scheme and asymptotically bounded by (Shannon) Entropy in a fashion that is similar to previous information and graph-theoretic approaches to chemical complexity [12, 13, 14, 15]. We conclude that Assembly Theory cannot offer a different or better explanation of selection or evolution than the connections already established [6].
@@ 2.1 The assembly Index is a compression algorithm of the LZ family
This limitation is attributable to the authors’ exclusive consideration of computer programs adhering to the form of ‘Template Program A’, as follows:
@@
Any approximation to $\mathbf{K}$ that accounts for identical repetitions, including all known lossless statistical compression algorithms, can achieve equivalent or superior results, as demonstrated in the Suplementary Information and [16]. This alignment with ‘Template Program A’ effectively highlights the association of Assembly Theory with well-established principles of compression and coding theory, thereby refuting the initial claim of its authors to present a unique methodology. Moreover, the authors’ suggestion that their index may be generalised as a universally applicable algorithm for any object (including text) [26]further underscores the disconnection of Assembly Theory and their authors’ suggestion to reinvent text compression (e.g. ZIP or GZIP) from the current state of the art in the field of statistical and non-statistical compression beyond LZW.
It is worth noting that although there may be minor variations in the implementation details of the assembly index and the assembly number of which the authors themselves have proposed significantly different versions, we demonstrate that they are all qualitatively and quantitatively equivalent to the LZ compression algorithms introduced in the 1970s, such as LZ77/LZ78. See Sup. Inf. A.
@@ 2.2 A compression algorithm is a mechanical procedure that corresponds to a physical process
The authors of AT argue that traditional data compression algorithms are an overly abstract process unsuited for modelling the construction (or assembly) of objects [26]. This view overlooks the practical and mechanistic nature of compression algorithms, particularly those in the LZ family. Since their introduction in 1977, LZ algorithms have been effectively used in detection, identification, clustering, and classification across various fields, including biology, chemistry, and medicine  [27, 15, 28], and to approximate algorithmic complexity [10, 29].
@@
On the contrary, one of the distinctive features of algorithmic information theory (AIT), algorithmic complexity being one of its indexes, is that its importance and pervasiveness in mathematics, theoretical computer science, and complexity science are, in fact, established because of the invariance of its results with respect to the model of computation, whether abstract or physically implemented. More specifically, such statements from the authors of AT [26] indicate a fundamental misunderstanding of equivalence classes closed under reductions (in this case, computability classes), which are basic and pivotal concepts in computer science, complexity theory, and mathematics in general. It is evidence in support of this point that none of the formal proofs in the Sup. Inf. section A made any mention of any Turing machine.
@@
In a published paper [4], the authors offered a proof of computability of their assembly index to distance themselves from Shannon Entropy and algorithmic (Kolmogorov) complexity. Their algorithm, categorised as Template Program A, is trivially computable and requires no proof of computability but also all other resource-bounded approximations to $\mathbf{K}$ are computable, including LZW that has been used for decades for this purpose [10]
@@ 2.3 Conflating object assembly process and causation directionality
The authors assume and present the sequential nature of their algorithm as an advantage [26]. Assembly Theory claims to be able to extract causal knowledge in the form of chains of cause and effect from its assembly pathways [26]. The authors’ central assumption is that each step of the basic Template Program A constitutes a cause-and-effect chain corresponding to how an object may have been physically assembled. This is no different from, and is indeed a restricted version of, a pushdown automaton capable of instantiating grammar compressor (which the authors rename as ‘assembly pathway’) and works exactly like an LZ algorithm.
@@
We have demonstrated that the assembly index requires a finite automaton to instantiate a restricted grammar (see Sup. Inf. A), a version of a specific-purpose machine, just as any other resource-bounded approximation to $\mathbf{K}$ would need to be instantiated (e.g. calculated, even by hand). In this case, the size of their formula or the size of the implementation of their computer program is the size of the special-purpose finite automaton, which is common and of fixed size for their calculations (which allows them to be discounted from the final length, just as it is from $\mathbf{K}$ or resource-bounded approximations).
@@ 2.4 Lack of control experiments and absence of supporting empirical data beyond current domain knowledge
Turing’s motivation was to explore what could be automated in the causal mechanisation of operations (see Fig. 1B), for example, stepwise by hand using paper and pencil. The shortest among all the computer programs of this type is an upper bound approximation to $\mathbf{K}$. In other words, $\mathbf{K}$ cannot be longer than the length of this diagram. The concept of pathway complexity and copy number instantiated by the assembly index, is a restricted version of statistical nature of a Turing machine or finite automaton. Moreover, a Turing machine is a synonym of an algorithm and vice versa. It is a mistake to think of a Turing machine as a physical object or an object with particular properties or features (such as a head or tape). All the algorithms in assembly theory are Turing machines and viceversa.
Figure 1: A: The authors of AT have suggested that $\mathbf{K}$ would be proven to be contained in Assembly TheoryT [26]. This Venn diagram shows how Assembly Theory (AT) is connected and subsumed within algorithmic complexity or ($\mathbf{K}$) and within the group of statistical compression as proven in this paper (see Sup. Inf. A). B: Causal transition graph of a Turing machine with number 3019 (in Wolfram’s enumeration scheme [33]) with an empty initial condition found by using a computable method (e.g. CTM, [34]) to explain how the block-patterned string 111000111000 was assembled step-by-step based on the principles of $\mathbf{K}$ describing the state, memory, and output of the process as a fully causal mechanistic explanation. A Turing machine is simply a procedural algorithm and any algorithm can be represented by a Turing machine. By definition, this is a mechanistic process, and as physical as anything else, not an ‘abstract’ or ‘unrealisable’ process.
2.5 A circular argument cannot unify physics and biology
Central to Assembly Theory (AT) and the public claims made through their University Press Releases [5] is the connection that the authors of AT want to make to selection and evolution, maintaining that AT unifies physics and biology and explains and quantifies selection and evolution, both Darwinian and abiotic [2]. To demonstrate this connection, by assuming selectivity in the combination of linear strings $(P)$, the authors compare two schemes for combining linear strings $Q$, one random versus a non-random selection. This experiment yields observations of differences between the two $(R)$, and the authors conclude that selectivity $(S)$ exists in molecular assembly and, therefore, that AT can explain it.
@@ 2.5 A circular argument cannot unify physics and biology
Yet, assuming selectivity in the combination of strings and proceeding to say that a selection algorithm is expected to differ from a random one makes for a circular argument. Furthermore, the conclusion $(S)$ reaffirms the initial assumption $(P)$, resulting in a circular argument.
@@
These circular arguments, lacking a foundation in chemistry, biology, or empirical evidence, lead the authors to propose the concept of “assembly time”. The suggestion is a time-scale separation between molecule production (assembly time) and discovery time, aiming to unify physics and biology. Time, however, has always been fundamental in evolutionary theory, and claims about the object history are the foundations of evolutionary theory. Assembly Theory revisits existing complexity science concepts described in the timeline provided in Fig. 2 but without acknowledgment or attribution.
Figure 2: A timeline of results in complexity science relevant to the claims and results from Assembly Theory. Assembly Theory renames several concepts such as dictionary trees as ‘assembly pathways’, heavily uses algorithmic probability in their reduction of combinatorial space arguments without attribution, and, as demonstrated, the assembly index and assembly number are equivalent to LZ lossless compression (proofs provided in the Sup. Inf. A)
@@ 3 Conclusions
The most popular examples used by the authors of Assembly Theory (AT) in their papers as illustrations of the way their algorithms work, such as ABRACADABRA and BANANA, are traditionally used to teach LZ77, LZ78, or LZW compression in Computer Science courses at university level. Dictionary trees have been used for pedagogic purposes in computer science for decades.
This would suggest some similarity between AT and compression algorithms, particularly those of a statistical nature based on recursion to previous states, copy counting, and the re-usage of patterns and repetitions.
In this paper, we have shown that the assembly index is a compression scheme belonging to the LZ family of compression algorithms, rather than only being similar. Diving into the details of their algorithms, we have revealed that they are equivalent to LZW and, therefore, to Shannon Entropy.
Our results demonstrate that an object with a low assembly index has high LZ compressibility (i.e., it is more compressible according to the LZ scheme) and, therefore, necessarily displays low entropy. In the opposite direction, an object with a high assembly index will have low LZ compressibility and, therefore, necessarily display high entropy. This means that Assembly Theory (AT), Shannon Entropy, and LZ compression algorithm are indistinguishable from each other, and the claim that those other approaches are incapable of dealing with their type of data, capture ‘structure’ or anything that could not have been already characterised by Shannon Entropy or LZW, as it has been done before, is inaccurate.
This paper also demonstrates that AT is subsumed into the theory and methods of Algorithmic Information Theory (AIT) as illustrated in Fig.1 1A and fully mathematically proven in the Sup. Inf. section A.
In addition, both theoretical concepts and empirical results in [1, 36, 4, 2] have been reported previously in earlier work in the context of AIT (from which algorithmic complexity is one of its indexes) for chemical processes in [15], and biology in [15, 37, 6, 38] but with comparison to other measures, including algorithms that go beyond statistical compression and connected to the concept of causal discovery.
@@
Algorithmic complexity has been shown to be profoundly connected to causality [39, 40, 24, 41], not only for strings, but also it has been applied to images [38], networks [7, 42, 43, 44], vectors, matrices, and tensors in multiple dimensions [45, 46, 47]; and chemical structures [15].
In contrast to statistical compression schemes of which AT is a particular case, exploring other resource-bounded computable approximations to algorithmic complexity [21, 24], that consider aspects other than statistics, such as identical repetitions, may have more discriminatory power.
This is because they may tell apart cases in which a high assembly index value with an expected low copy number (or frequency of occurrence in the ensemble), low LZ compressibility and high entropy, would indicate that an object is more causally disconnected, independent, or statistically random, but actually is causally dependent due to being more compressible by exploiting, for example, short rewriting rules describing the object.
@@
One positive aspect of Assembly Theory is that it offers a graph-like representation approach (albeit more convoluted than necessary) to approximate LZ compression in the specific context of molecular complexity, potentially making it more accessible to a broader and less technical community. However, it is inaccurate and a bad practice to ignore, omit or imply the absence of previous work (e.g. [15, 6, 48]), belittle and devalue the theories and ideas upon which their index is entirely dependable and connected with (Shannon Entropy, compression, and algorithmic Kolmogorov complexity), and make disproportionate public claims related to the contributions of AT (even calling it a theory that ‘unifies biology and physics’ and a ‘theory of everything’ by the media and the authors).
We agree with the authors on the importance of understanding the origin and mechanisms of the emergence of an open-ended generation of novelty [49, 41]. However, these efforts by the authors are undermined by hyperbolic claims.
We argue that the results and claims made by AT, in fact, highlight a working hypothesis that information and computation underpin the concepts necessary to explain the processes and building blocks of physical and living systems, concepts which the methods and frameworks of AT have been heavily inspired upon [50, 51].
@@ A.3.1 Assembly index defines an encoding scheme
    The following results show that the assembly index $c_{\Gamma}\left(y\right)$ is a parameter that can be employed to encode an object $y$. Theorem 5 demonstrates that any arbitrarily chosen encoding of the set of vertices (in any arbitrary ordering) of a minimum rooted assembly space is sufficient to encode the highest-assembly-index object. In other words, the sheer collection of objects in a minimum rooted assembly subspace carries enough information to allow one to always pick the object with the highest assembly index value, even though the ordering of the objects in this collection can be completely arbitrary in the first place. This demonstrates that Assembly Theory is one of the possible methods to encode objects.